I really need your help, I tried this and keep coming up with the wrong answer.-------> solve the following equation for x...x+2/x+4=x-4/x+5... the answer I got was -26/7. thanks
Answer by linh117(31)

x+(2)/(x)+4=x-(4)/(x)+5
Find the LCD (least common denominator) of x+(2)/(x)+4+x-(4)/(x)+5.
Least common denominator: x
Multiply each term in the equation by x in order to remove all the denominators from the equation.
x*x+(2)/(x)*x+4*x=x*x-(4)/(x)*x+5*x
Simplify the left-hand side of the equation by canceling the common terms.
x^(2)+4x+2=x*x-(4)/(x)*x+5*x
Simplify the right-hand side of the equation by simplifying each term.
x^(2)+4x+2=x^(2)+5x-4
Move all terms not containing x to the right-hand side of the equation.
x^(2)+4x+2-x^(2)-5x=-4
Since x^(2) and -x^(2) are like terms, add -x^(2) to x^(2) to get 0.
0+4x+2-5x=-4
Since 4x and -5x are like terms, add -5x to 4x to get -x.
-x+2=-4
Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides.
-x=-2-4
Subtract 4 from -2 to get -6.
-x=-6
Multiply each term in the equation by -1.
-x*-1=-6*-1
Multiply -x by -1 to get x.
x=-6*-1
Multiply -6 by -1 to get 6.
x=6